OpenStudy (anonymous):

Solve the following equation: x/x-5=2/x-3+4/(x-3)(x-5) (find x)

OpenStudy (anonymous):

can you please write the equation by using the equation feature

OpenStudy (anonymous):

hello are you there?

OpenStudy (anonymous):

\[\frac{ x }{ x-5 }=\frac{ 2 }{ x-3 }+\frac{ 4 }{ (x-3)(x-5) }\]

OpenStudy (anonymous):

these are the steps

OpenStudy (anonymous):

answer is x=2 or x=3

OpenStudy (anonymous):

sorry! I didn't see you responded

OpenStudy (anonymous):

thank you so much! it helped